# Privacy-Preserving Federated Learning via Homomorphic Adversarial Networks

### v2

@@ Abstract
Privacy-preserving federated learning (PPFL) aims to train a global model for multiple clients while maintaining their data privacy. However, current PPFL protocols exhibit one or more of the following insufficiencies: considerable degradation in accuracy, the requirement for sharing keys, and cooperation during the key generation or decryption processes. As a mitigation, we develop the first protocol that utilizes neural networks to implement PPFL, as well as incorporating an Aggregatable Hybrid Encryption scheme tailored to the needs of PPFL. We name these networks as Homomorphic Adversarial Networks (HANs) which demonstrate that neural networks are capable of performing tasks similar to multi-key homomorphic encryption (MK-HE) while solving the problems of key distribution and collaborative decryption. Our experiments show that HANs are robust against privacy attacks. Compared with non-private federated learning, experiments conducted on multiple datasets demonstrate that HANs exhibit a negligible accuracy loss (at most 1.35%). Compared to traditional MK-HE schemes, HANs increase encryption aggregation speed by 6,075 times while incurring a 29.2$\times$ increase in communication overhead.
@@ Introduction
    We pioneer the use of neural networks to emulate MK-HE algorithms and optimize encryption and aggregation in FL through our AHE scheme. AHE provides irreversible ciphertext and a private key encryption mechanism for individual $clients$, aiming to offer new insights into privacy protection research in neural network-based cryptography and FL protocols.
@@
    HANs with AHE balances privacy, performance, and efficiency. It eliminates collaborative decryption, key sharing, and collective key generation among clients. The framework enables OTP usage with minimal cost and ensures privacy even if $N$-$2$ $Clients$ and $Servers$ collude.
@@
    Empirical evaluations across multiple datasets demonstrate the efficiency and effectiveness of HANs with AHE. Compared to non-private FL, it shows at most only 1.35% accuracy loss. HANs increase encryption aggregation speed by 6,075 times compared to traditional MK-HE schemes while incurring a 29.2$\times$ increase in communication overhead.
@@
In order to prevent collusion attacks, MK-HE allows multiple parties to utilize distinct keys for encryption. The decryption process necessitates the collaborative involvement of all parties. Ma et al. [23] introduced a PPFL framework based on xMK-CKKS, demonstrating resilience against collusion involving fewer than $N$-$1$ participant devices and the $Server$.
@@
The security model involves a $Server$, multiple $Clients$, and an $attacker$. For ease of illustration, we assume three $Clients$ in this scenario, each aiming to protect the privacy of its local dataset $D_{i}$ from both the $Server$ and other $Clients$.
@@
    $(pk,sk)\leftarrow\text{KeyGen}(\kappa)$. Generates a public key $pk$ and private keys $sk$ =$=\{sk_{A},sk_{B}\}$.
@@
[…]
@@
Pseudo $N$-$1$ Collusion Attacks
In addition to attacks and challenges targeting the model’s inherent encryption capabilities, the unique characteristics of HANs may lead to two types of pseudo $N$-$1$ collusion attacks. These attacks attempt to overcome the limitations of traditional $N$-$2$ collusion attacks by leveraging additional information to achieve an effect approximating $N$-$1$ collusion. However, due to the PPU mechanism, their effectiveness remains significantly limited. The basic ideas behind these two attacks are:
@@
    Pseudo $N$-$1$ Collusion Attack based on Original Model (PCAOM): In this attack, the adversary uses another trusted client’s original model to substitute for that client’s private model. This is a KMA where the attacker attempts to simulate collusion among $N$-$1$ clients by using the publicly available original model, while in reality only $N$-$2$ clients are colluding.
@@
    Pseudo $N$-$1$ Collusion Attack based on Public Dataset (PCAPD):
    This attack utilizes the noisy public dataset’s information generated during the PPU process. The attacker uses this public data to approximate the behavior of another trusted client, thereby achieving an effect similar to $N$-$1$ collusion. This is an enhanced version of a COA.
@@
Resistance to Pseudo $N$-$1$ Collusion Attacks
In evaluating the security of HANs, we conducted experiments on pseudo $N$-$1$ collusion attacks. Table 4 presents the results of the PCAOM and PCAPD.
These attack methods attempt to simulate the effect of $N$-$1$ collusion attacks, but their effectiveness is significantly limited due to the PPU mechanism. The experimental results show that after implementing PPU, PCAOM has a MAD of 0.31067, while PCAPD has a MAD of 0.30340. Notably, before the implementation of PPU, the result of PCAOM was equivalent to the Average value of HANs, indicating that the PPU mechanism effectively enhanced the system’s security.
@@
Table 5: Performance metrics of HANs for various batch sizes
Batch Size Batch Encryption Time Batch Aggregation Time Key Generation Time
100,000 0.020451s 0.018023s 0.000034s
200,000 0.035654s 0.035722s 0.000029s
300,000 0.053236s 0.053374s 0.000027s
@@
Our results demonstrate that HANs significantly outperforms SecFed in terms of computational efficiency. For a batch of 3,000 ciphertexts, HANs completes encryption and aggregation in just 0.00107 seconds, compared to SecFed’s 6.5 seconds. This represents a remarkable 6,075$\times$ speedup, highlighting the exceptional computational efficiency of our approach. The significant improvement in computational efficiency comes at a cost, specifically a 29.2$\times$ increase in communication overhead compared to SecFed.
@@ Conclusion
This work introduces Homomorphic Adversarial Networks (HANs) with Aggregatable Hybrid Encryption for Privacy-Preserving Federated Learning (PPFL). HANs leverage neural networks to emulate multi-key homomorphic encryption, offering a novel approach that balances privacy, performance, and efficiency. Our method enables independent key generation and aggregation without collaborative decryption, while resisting $N$-$2$ client collusion. The innovative Privacy-Preserving Update mechanism enhances security through private model updates, effectively mitigating potential vulnerabilities in the initial public model. Experimental results demonstrate HANs’ ability to maintain model accuracy within 1.35% of non-private federated learning. HANs also significantly outperform traditional multi-key homomorphic encryption schemes, achieving a 6,075$\times$ increase in computational efficiency. The introduction of these neural network-based protocols not only improves the practical implementation of PPFL but also opens new research directions in federated learning privacy protocols and neural network-based cryptography.
@@ Appendix A Problem Overview
Suppose that we have $m$ separate $Clients$. Each $Client$ is represented by $C_{i}$, where $i\in[1,m]$ and $Client$ $C_{i}$ has a local training dataset $\mathcal{D}_{i}$. In each step, there are three sub-steps:
@@
    Broadcast. $Server$ broadcasts the current global model parameters $w^{t-1}$ to each $Client$ $C_{i}$, where $t$ represents the index of the current iteration round.
@@
    Local training. Each $Client$ $C_{i}$ receives global model parameters $w^{t-1}$ and using local training datasets $\mathcal{D}_{i}$ to obtain the new local model parameters $w^{t}_{i}$ in parallel and sends the local model parameter $w^{t}_{i}$ back to the $Server$. During the updating step, the $Client$ typically employs stochastic gradient descent for local epochs. In scenarios where communication cost is not a primary concern, setting local epochs to 1 can be an effective approach [25].
@@
    Aggregation. The $Server$ receives all model parameters {$w^{t}_{1}$,$w^{t}_{2}$$\dots$,$w^{t}_{m}$} from the $Client$ and aggregates them into global parameters $w^{t}$ by averaging them.
@@ Appendix B HANs Training Design: Loss Function Formulation and Rationale
The preceding discussion focuses on optimization methods for attackers. As for encryption, intuitively, it can be viewed as a hindrance to the attackers’ goals. These goals can be achieved by maximizing the loss for each $client$ in the $Clients$ list, thus realizing the optimal encryption model:
@@ Appendix C HANs Training Implementation: A Multi-stage Optimization Process
Appendix D Pseudo $N$-$1$ Collusion Attacks
PPFL scenarios based on multi-key homomorphic encryption typically can only resist $N$-$2$ collusion attacks. This is because if only one client remains honest and trustworthy, colluding clients can easily obtain that client’s real data by subtracting their uploaded data from the aggregated model gradients. This scenario is not one that multi-key homomorphic encryption is designed to defend against, and our method is no exception. However, due to the unique characteristics of HANs, attackers may potentially employ two types of pseudo $N$-$1$ collusion attacks. We will now formally define these two attacks.
Pseudo $N$-$1$ Collusion Attack Based on the Original Model (PCAOM)
PCAOM is a KMA. Let $clients={client_{1},\ldots,client_{N}}$ be a set of $N$ $clients$ in a FL system using HANs. This attack can be described in the following steps:
@@ Appendix D Pseudo $N$-$1$ Collusion Attacks
    • The $N$-$2$ colluding attackers acquire the latest aggregation model $Aggregation()$.
@@
Pseudo $N$-$1$ Collusion Attack Based on Public Dataset (PCAPD)
@@
    • The $N$-$2$ colluding attackers acquire the latest aggregation model.
@@
      • A noisy plaintext $m_{B}^{pub}=m_{B}+noise$, where $noise$ is unknown to the attacker
@@ Appendix E PPU Mechanism
  • Diminishing the correlation between the public dataset and the $client$’s encryption model.
@@ Appendix G Security Discussion of HANs
Guessing $\text{Guessing Difference}=|m_{\text{guess}}-m_{\text{real}}|$ (3)
@@
To achieve the attack objective, attackers will employ various strategies to implement attacks. Although we cannot enumerate all possible attack methods, we can discuss the difficulty of attacks and the security of HANs by analyzing the information potentially exposed to attackers. Assuming HANs is used by at least two honest $clients$, $client_{A}$ and $client_{B}$, we will discuss the scenario where an attacker attempts to obtain information from $client_{A}$.
@@
    HANs usage phase: $\{pk_{i},c_{i}\}_{i=1}^{N},m_{\text{Agg}}$ where $pk_{i}$ and $c_{i}$ represent the public key and ciphertext of the $i$-th $client$, respectively, and $m_{\text{Agg}}$ represents the aggregated value of the ciphertexts.
@@
However, according to our security protocol, honest $clients$ must add noise to their public datasets, making it difficult to satisfy the above conditions and significantly increasing the difficulty of attacks.
@@
8:         $l\leftarrow\text{RandomInteger}(1,|D_{j}^{pub}|)$triangleright Randomly select an index from the current client’s dataset
@@
1:Number of clients $N$, maximum iterations $maxIterations$, pre-trained HANs model (including encryption model $\theta^{enc}$ and aggregation model $\theta^{agg}$), private dataset size $privateDataSize$, public dataset size $publicDataSize$, noise scale $\sigma$
@@
5:     $noise\leftarrow\text{generateGaussianNoise}(\sigma)$triangleright Generate initial Gaussian noise
6:     $x_{i}^{noisy}\leftarrow x_{i}^{init}+noise$triangleright Add initial noise to plaintext
@@
8:     $D_{i}^{pub}\leftarrow{x_{i}^{noisy},(sk_{i1}^{init}+sk_{i2}^{init}),c_{i}^{init}}$triangleright Initialize public dataset with noisy plaintext
@@
10:for $t=1$ to $maxIterations$ do
11:     for $i=1$ to $N$ do
12:         $D_{i}^{others}\leftarrow\bigcup_{j=1,j\neq i}^{N}D_{j}^{pub}$triangleright Collect all other clients’ public datasets
13:         $x_{i}^{own},sk_{i1},sk_{i2}\leftarrow\text{generateRandomData}(privateDataSize)$triangleright Generate private data
@@
15:         $\theta_{i}^{enc,new},\theta_{i}^{agg,new}\leftarrow\text{Optimize}(\theta_{i}^{enc},\theta_{i}^{agg},T_{i})$triangleright Update both models
@@
17:         $c_{i}^{new}\leftarrow\text{Encrypt}(x_{i}^{new},sk_{i1}^{new},sk_{i2}^{new},\theta_{i}^{enc,new})$triangleright Encrypt using updated encryption
18:         $noise\leftarrow\text{generateGaussianNoise}(\sigma)$triangleright Generate Gaussian noise with increased scale
19:         $x_{i}^{noisy}\leftarrow$ Agg$(x_{i}^{new}+noise)$triangleright Apply aggregation function on noisy plaintext
20:         $D_{i}^{pub}\leftarrow{x_{i}^{noisy},(sk_{i1}^{new}+sk_{i2}^{new}),c_{i}^{new}}$triangleright Update public dataset
21:         $\theta_{i}^{enc}\leftarrow\theta_{i}^{enc,new}$triangleright Update encryption model
@@
1:Number of clients $N$, rounds per client $roundsPerClient$, HANs model after CPPU (including client-specific encryption models $\{\theta_{i}^{enc}\}_{i=1}^{N}$ and aggregation model $\theta^{agg}$), private dataset size $privateDataSize$, final public datasets from original CPPU $\{D_{i}^{pub}\}_{i=1}^{N}$
@@
3:for $i=1$ to $N$ do $\triangleright$ This loop can be executed in parallel for each client
4:     $D_{i}^{others}\leftarrow\bigcup_{j=1,j\neq i}^{N}D_{j}^{pub}$triangleright Collect all other clients’ public datasets
5:     for $r=1$ to $roundsPerClient$ do
6:         $x_{i}^{own},sk_{i1},sk_{i2}\leftarrow\text{generateRandomData}(privateDataSize)$triangleright Generate private data
@@
8:         $\theta_{i}^{enc}\leftarrow\text{Optimize}(\theta_{i}^{enc},\theta^{agg},T_{i})$triangleright Update enc model
@@
11:return $\{\theta_{i}^{enc}\}_{i=1}^{N}$triangleright Return final updated encryption models
@@
4:     $x^{\prime}_{1}\leftarrow\mathcal{N}(0,1)$, $y^{\prime}_{1}\leftarrow\mathcal{N}(0,1)$triangleright Initialize dummy inputs and labels
@@
6:         $\nabla W^{\prime}_{i}\leftarrow\frac{\partial\ell(F(x^{\prime}_{i},W_{t}),y^{\prime}_{i})}{\partial W_{t}}$triangleright Compute dummy gradients
7:         $D_{i}\leftarrow||\nabla W^{\prime}_{i}-\nabla W||^{2}$
8:         $x^{\prime}_{i+1}\leftarrow x^{\prime}_{i}-\eta\nabla_{x^{\prime}_{i}}D_{i}$, $y^{\prime}_{i+1}\leftarrow y^{\prime}_{i}-\eta\nabla_{y^{\prime}_{i}}D_{i}$triangleright Update data to match gradients
@@
1:$F(x;W)$: Differentiable machine learning model; $W$: model parameters; $Enc$: encryption function; $sk_{1},sk_{2}$: secret keys; $Attack_{1},Attack_{2}$: adversary’s attack models
@@
3:procedure DLG-HANs($F$, $W$, $Enc$, $sk_{1}$, $sk_{2}$, $Attack_{1}$, $Attack_{2}$)
@@
5:     $c\leftarrow Enc(sk_{1},sk_{2},\nabla W)$triangleright Client encrypts gradients
6:     $pk\leftarrow sk_{1}+sk_{2}$triangleright Public key is sum of secret keys
7:     $\nabla W_{1}\leftarrow Attack_{1}(pk,c)$triangleright Attack using public key and ciphertext
8:     $\nabla W_{2}\leftarrow Attack_{2}(c)$triangleright Attack using only ciphertext
9:     $(x^{\prime}_{1},y^{\prime}_{1})\leftarrow\text{DLG}(F,W,\nabla W_{1})$triangleright Apply DLG on first attack result
10:     $(x^{\prime}_{2},y^{\prime}_{2})\leftarrow\text{DLG}(F,W,\nabla W_{2})$triangleright Apply DLG on second attack result

### v3

# {textblock}

@@ Abstract
[…]
@@ Introduction
    We pioneer the use of neural networks to emulate MK-HE algorithms and optimize encryption and aggregation in FL through our AHE scheme. AHE provides irreversible ciphertext and a private key encryption mechanism for individual clientsclients, aiming to offer new insights into privacy protection research in neural network-based cryptography and FL protocols.
@@
    HANs with AHE balances privacy, performance, and efficiency. It eliminates collaborative decryption, key sharing, and collective key generation among clients. The framework enables OTP usage with minimal cost and ensures privacy even if NN-22 ClientsClients and $Servers$ collude.
@@
[…]
@@
In order to prevent collusion attacks, MK-HE allows multiple parties to utilize distinct keys for encryption. The decryption process necessitates the collaborative involvement of all parties. Ma et al. [23] introduced a PPFL framework based on xMK-CKKS, demonstrating resilience against collusion involving fewer than NN-11 participant devices and the $Server$.
@@
The security model involves a $Server$, multiple $Clients$, and an attackerattacker. For ease of illustration, we assume three $Clients$ in this scenario, each aiming to protect the privacy of its local dataset $D_{i}$ from both the $Server$ and other $Clients$.
@@
    $(pk,sk)\leftarrow\text{KeyGen}(\kappa)$. Generates a public key $pk$ and private keys sksk $\{sk_{A},sk_{B}\}$.
@@
$\textit{Attack successful.}\Leftrightarrow D^{attack1}_{i}\simeq D_{i}\ \cup\ D^{attack2}_{i}\simeq D_{i}$
@@
Pseudo NN-11 Collusion Attacks
In addition to attacks and challenges targeting the model’s inherent encryption capabilities, the unique characteristics of HANs may lead to two types of pseudo NN-11 collusion attacks. These attacks attempt to overcome the limitations of traditional NN-22 collusion attacks by leveraging additional information to achieve an effect approximating NN-11 collusion. However, due to the PPU mechanism, their effectiveness remains significantly limited. The basic ideas behind these two attacks are:
@@
    Pseudo NN-11 Collusion Attack based on Original Model (PCAOM): In this attack, the adversary uses another trusted client’s original model to substitute for that client’s private model. This is a KMA where the attacker attempts to simulate collusion among NN-11 clients by using the publicly available original model, while in reality only NN-22 clients are colluding.
@@
    Pseudo NN-11 Collusion Attack based on Public Dataset (PCAPD):
    This attack utilizes the noisy public dataset’s information generated during the PPU process. The attacker uses this public data to approximate the behavior of another trusted client, thereby achieving an effect similar to NN-11 collusion. This is an enhanced version of a COA.
@@
Resistance to Pseudo NN-11 Collusion Attacks
In evaluating the security of HANs, we conducted experiments on pseudo NN-11 collusion attacks. Table 4 presents the results of the PCAOM and PCAPD.
These attack methods attempt to simulate the effect of NN-11 collusion attacks, but their effectiveness is significantly limited due to the PPU mechanism. The experimental results show that after implementing PPU, PCAOM has a MAD of 0.31067, while PCAPD has a MAD of 0.30340. Notably, before the implementation of PPU, the result of PCAOM was equivalent to the Average value of HANs, indicating that the PPU mechanism effectively enhanced the system’s security.
@@
Table 5: Performance metrics of HANs (results based on 1000 experiments)
Batch Size Encryption Time Aggregation Time Key Generation Time
100,000 0.019554s (±0.001629) 0.017444s (±0.000108) 0.000028s (±0.000008)
200,000 0.035329s (±0.000027) 0.035380s (±0.000013) 0.000029s (±0.000008)
300,000 0.053281s (±0.003922) 0.053462s (±0.004456) 0.000031s (±0.000012)
@@
[…]
@@ Conclusion
[…]
@@ Appendix A Problem Overview
Suppose that we have $m$ separate $Clients$. Each $Client$ is represented by $C_{i}$, where $i\in[1,m]$ and ClientClient $C_{i}$ has a local training dataset $\mathcal{D}_{i}$. In each step, there are three sub-steps:
@@
    Broadcast. $Server$ broadcasts the current global model parameters $w^{t-1}$ to each ClientClient $C_{i}$, where $t$ represents the index of the current iteration round.
@@
    Local training. Each ClientClient $C_{i}$ receives global model parameters $w^{t-1}$ and using local training datasets $\mathcal{D}_{i}$ to obtain the new local model parameters $w^{t}_{i}$ in parallel and sends the local model parameter $w^{t}_{i}$ back to the $Server$. During the updating step, the $Client$ typically employs stochastic gradient descent for local epochs. In scenarios where communication cost is not a primary concern, setting local epochs to 1 can be an effective approach [25].
@@
Definition of $\delta$-accuracy loss. Suppose that $\mathcal$$\bar${M}} is a deep learning model training on datasets $\mathcal{D}$, where $\mathcal{D}=\mathcal{D}_{i}\cup\mathcal{D}_{2}\cup\dots\cup\mathcal{D}_{m}$. We use $\boldsymbol{\bar{f}}$ to denote the accuracy of model $\mathcal$$\bar${M}}. For FL, $\mathcal$$\hat${M}} denotes the model after all train rounds, and its corresponding accuracy is $\boldsymbol{\hat{f}}$. We say that it is $\delta$-accuracylossaccuracy\ loss, if it satisfies $\bar{f}-\hat{f}<\delta$.
@@ Appendix B HANs Training Design: Loss Function Formulation and Rationale
The preceding discussion focuses on optimization methods for attackers. As for encryption, intuitively, it can be viewed as a hindrance to the attackers’ goals. These goals can be achieved by maximizing the loss for each clientclient in the $Clients$ list, thus realizing the optimal encryption model:
@@ Appendix C HANs Training Implementation: A Multi-stage Optimization Process
Appendix D Pseudo NN-11 Collusion Attacks
PPFL scenarios based on multi-key homomorphic encryption typically can only resist NN-22 collusion attacks. This is because if only one client remains honest and trustworthy, colluding clients can easily obtain that client’s real data by subtracting their uploaded data from the aggregated model gradients. This scenario is not one that multi-key homomorphic encryption is designed to defend against, and our method is no exception. However, due to the unique characteristics of HANs, attackers may potentially employ two types of pseudo NN-11 collusion attacks. We will now formally define these two attacks.
Pseudo NN-11 Collusion Attack Based on the Original Model (PCAOM)
PCAOM is a KMA. Let $clients={client_{1},\ldots,client_{N}}$ be a set of NN clientsclients in a FL system using HANs. This attack can be described in the following steps:
@@ Appendix D Pseudo NN-11 Collusion Attacks
    • The NN-22 colluding attackers acquire the latest aggregation model $Aggregation()$.
@@
Pseudo NN-11 Collusion Attack Based on Public Dataset (PCAPD)
@@
    • The NN-22 colluding attackers acquire the latest aggregation model.
@@
      • A noisy plaintext $m_{B}^{pub}=m_{B}+noise$, where noisenoise is unknown to the attacker
@@ Appendix E PPU Mechanism
  • Diminishing the correlation between the public dataset and the clientclient’s encryption model.
@@ Appendix F Detailed Explanation of DLG Attack
$\textit{Attack successful.}\Leftrightarrow x^{\prime}\simeq x\ \cap\ y^{\prime}\simeq y$
@@ Appendix G Security Discussion of HANs
$\text{Guessing Difference}=|m_{\text{guess}}-m_{\text{real}}|$ (3)
@@
To achieve the attack objective, attackers will employ various strategies to implement attacks. Although we cannot enumerate all possible attack methods, we can discuss the difficulty of attacks and the security of HANs by analyzing the information potentially exposed to attackers. Assuming HANs is used by at least two honest clientsclients, $client_{A}$ and $client_{B}$, we will discuss the scenario where an attacker attempts to obtain information from $client_{A}$.
@@
    HANs usage phase: $\{pk_{i},c_{i}\}_{i=1}^{N},m_{\text{Agg}}$ where $pk_{i}$ and $c_{i}$ represent the public key and ciphertext of the $i$-th clientclient, respectively, and $m_{\text{Agg}}$ represents the aggregated value of the ciphertexts.
@@
However, according to our security protocol, honest clientsclients must add noise to their public datasets, making it difficult to satisfy the above conditions and significantly increasing the difficulty of attacks.
@@
8:   $l\leftarrow\text{RandomInteger}(1,|D_{j}^{pub}|)$ $\triangleright$ Randomly select an index from the current client’s dataset
@@
1:Number of clients NN, maximum iterations maxIterationsmaxIterations, pre-trained HANs model (including encryption model $\theta^{enc}$ and aggregation model $\theta^{agg}$), private dataset size privateDataSizeprivateDataSize, public dataset size publicDataSizepublicDataSize, noise scale $\sigma$
@@
5:  $noise\leftarrow\text{generateGaussianNoise}(\sigma)$ $\triangleright$ Generate initial Gaussian noise
6:  $x_{i}^{noisy}\leftarrow x_{i}^{init}+noise$ $\triangleright$ Add initial noise to plaintext
@@
8:  $D_{i}^{pub}\leftarrow{x_{i}^{noisy},(sk_{i1}^{init}+sk_{i2}^{init}),c_{i}^{init}}$ $\triangleright$ Initialize public dataset with noisy plaintext
@@
10:for $t=1$ to maxIterationsmaxIterations do
11:  for $i=1$ to NN do
12:   $D_{i}^{others}\leftarrow\bigcup_{j=1,j\neq i}^{N}D_{j}^{pub}$ $\triangleright$ Collect all other clients’ public datasets
13:   $x_{i}^{own},sk_{i1},sk_{i2}\leftarrow\text{generateRandomData}(privateDataSize)$ $\triangleright$ Generate private data
@@
15:   $\theta_{i}^{enc,new},\theta_{i}^{agg,new}\leftarrow\text{Optimize}(\theta_{i}^{enc},\theta_{i}^{agg},T_{i})$ $\triangleright$ Update both models
@@
17:   $c_{i}^{new}\leftarrow\text{Encrypt}(x_{i}^{new},sk_{i1}^{new},sk_{i2}^{new},\theta_{i}^{enc,new})$ $\triangleright$ Encrypt using updated encryption
18:   $noise\leftarrow\text{generateGaussianNoise}(\sigma)$ $\triangleright$ Generate Gaussian noise with increased scale
[…]
20:   $D_{i}^{pub}\leftarrow{x_{i}^{noisy},(sk_{i1}^{new}+sk_{i2}^{new}),c_{i}^{new}}$ $\triangleright$ Update public dataset
21:   $\theta_{i}^{enc}\leftarrow\theta_{i}^{enc,new}$ $\triangleright$ Update encryption model
@@
1:Number of clients NN, rounds per client roundsPerClientroundsPerClient, HANs model after CPPU (including client-specific encryption models $\{\theta_{i}^{enc}\}_{i=1}^{N}$ and aggregation model $\theta^{agg}$), private dataset size privateDataSizeprivateDataSize, final public datasets from original CPPU $\{D_{i}^{pub}\}_{i=1}^{N}$
@@
3:for $i=1$ to NN do $\triangleright$ This loop can be executed in parallel for each client
4:  $D_{i}^{others}\leftarrow\bigcup_{j=1,j\neq i}^{N}D_{j}^{pub}$ $\triangleright$ Collect all other clients’ public datasets
5:  for $r=1$ to roundsPerClientroundsPerClient do
6:   $x_{i}^{own},sk_{i1},sk_{i2}\leftarrow\text{generateRandomData}(privateDataSize)$ $\triangleright$ Generate private data
@@
8:   $\theta_{i}^{enc}\leftarrow\text{Optimize}(\theta_{i}^{enc},\theta^{agg},T_{i})$ $\triangleright$ Update enc model
@@
11:return $\{\theta_{i}^{enc}\}_{i=1}^{N}$ $\triangleright$ Return final updated encryption models
@@
4:  $x^{\prime}_{1}\leftarrow\mathcal{N}(0,1)$, $y^{\prime}_{1}\leftarrow\mathcal{N}(0,1)$ $\triangleright$ Initialize dummy inputs and labels
@@
6:   $\nabla W^{\prime}_{i}\leftarrow\frac{\partial\ell(F(x^{\prime}_{i},W_{t}),y^{\prime}_{i})}{\partial W_{t}}$ $\triangleright$ Compute dummy gradients
[…]
8:   $x^{\prime}_{i+1}\leftarrow x^{\prime}_{i}-\eta\nabla_{x^{\prime}_{i}}D_{i}$, $y^{\prime}_{i+1}\leftarrow y^{\prime}_{i}-\eta\nabla_{y^{\prime}_{i}}D_{i}$ $\triangleright$ Update data to match gradients
@@
1:$F(x;W)$: Differentiable machine learning model; $W$: model parameters; EncEnc: encryption function; $sk_{1},sk_{2}$: secret keys; $Attack_{1},Attack_{2}$: adversary’s attack models
@@
3:procedure DLG-HANs($F$, $W$, EncEnc, $sk_{1}$, $sk_{2}$, $Attack_{1}$, $Attack_{2}$)
@@
5:  $c\leftarrow Enc(sk_{1},sk_{2},\nabla W)$ $\triangleright$ Client encrypts gradients
6:  $pk\leftarrow sk_{1}+sk_{2}$ $\triangleright$ Public key is sum of secret keys
7:  $\nabla W_{1}\leftarrow Attack_{1}(pk,c)$ $\triangleright$ Attack using public key and ciphertext
8:  $\nabla W_{2}\leftarrow Attack_{2}(c)$ $\triangleright$ Attack using only ciphertext
9:  $(x^{\prime}_{1},y^{\prime}_{1})\leftarrow\text{DLG}(F,W,\nabla W_{1})$ $\triangleright$ Apply DLG on first attack result
10:  $(x^{\prime}_{2},y^{\prime}_{2})\leftarrow\text{DLG}(F,W,\nabla W_{2})$ $\triangleright$ Apply DLG on second attack result